# PersoBench: Benchmarking Personalized Response Generation in Large Language Models

### v1

@@ Abstract
While large language models (LLMs) have exhibited impressive conversational capabilities, their proficiency in delivering personalized responses remains unclear. Although recent benchmarks automatically evaluate persona consistency in role-playing contexts using LLM-based judgment, the evaluation of personalization in response generation remains underexplored. To address this gap, we present a new benchmark, PersoBench, to evaluate the personalization ability of LLMs in persona-aware dialogue generation within a zero-shot setting. We assess the performance of three open-source and three closed-source LLMs using well-known datasets and a range of metrics. Our analysis, conducted on three well-known persona-aware datasets, evaluates multiple dimensions of response quality, including fluency, diversity, coherence, and personalization, across both standard and chain-of-thought prompting methods. Our findings reveal that while LLMs excel at generating fluent and diverse responses, they are far from satisfactory in delivering personalized and coherent responses considering both the conversation context and the provided personas. Our benchmark implementation is available at https://github.com/salehafzoon/PersoBench.
Keywords Personalized Response Generation  $\cdot$ Conversational AI, $\cdot$ Natural Language Processing $\cdot$ Large Language Models $\cdot$ Benchmarking
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized NLP, excelling in human-like text generation across domains and becoming central to dialogue systems. However, evaluating their ability to generate personalized responses that enhance user engagement is crucial, especially in applications like customer service, where tailored interactions boost satisfaction [1].
While recent benchmarks such as RPBench-Auto [2], TIMECHARA [3] and RoleLLM [4] have been introduced in the role-playing domain to assess LLMs’ adherence to predefined characters or roles in character-based, scene-based, and temporal setups, there is still no dedicated benchmark for automatic personalized response generation of LLMs in the literature. Further, existing benchmarks also suffer from biases in their evaluations due to the use of large LLMs as judges, and limited experimental sizes constrain them.
To fill this gap, we introduce PersoBench, a benchmark for response personalization, to assess the strengths and limitations of current LLMs in generating personalized responses. To the best of our knowledge, no prior work has introduced a comprehensive benchmark specifically focused on evaluating response personalization in LLMs. Using comprehensive datasets and a diverse set of established metrics, including fluency, diversity, and coherence, we ensure a robust evaluation of various aspects of response generation, drawing on insights from a recent survey in the field [1]. More specifically, in line with this objective of the mentioned context, we aim to answer the following research questions:
@@
    Can LLMs generate fluent responses?
@@
    Can LLMs generate responses that maintain coherence with the given context and persona?
@@
    Can LLMs generate personalized responses?
To answer the above questions, we benchmarked three open-source and three closed-source LLMs across three popular persona-aware conversational datasets. Their performance is evaluated in a zero-shot setting, considering both standard and Chain of Thought (COT) prompting configurations. Our evaluation goes beyond assessing overall response quality and contextualization, employing eight established metrics in conversational AI to gauge performance across fluency, diversity, coherence, and personalization dimensions. This work contributes by (1) offering a multi-aspect evaluation of LLM personalization in zero-shot settings, (2) investigating the influence of COT prompting on response personalization, and (3) assessing response time and instructability (the ability to adhere to task instructions) based on LLMs’ adherence to task instructions.
@@ 2 Related Work
2.1 Existing Studies on LLM Benchmarking
LLM benchmarking has explored multiple aspects of their performance, including evaluations of factuality, ethical behaviour, task controllability, domain-specific applications, and cost efficiency. Besides their evaluation context, these benchmarks can be categorized based on the level of human involvement, which affects the degree of automation, and by their evaluation methods—whether they use explicit metrics for objective assessments or rely on LLMs as evaluators to judge model outputs.
Benchmarks relying on human-in-the-loop approaches require human involvement for tasks like error detection, ethical evaluation, and annotation. FELM [5], for example, uses human annotators to evaluate the factual accuracy of LLM outputs across domains like math, science, and reasoning. It focuses on granular annotations to capture subtle errors, providing detailed error types and reference links. TrustGPT [6] integrates human oversight for evaluating ethical concerns such as bias and value alignment, using tools like the PERSPECTIVE API [7] for toxicity scores while relying on human judgment for subjective issues.
On the other hand, some studies adopt the LLM-as-a-Judge approach for automated evaluation. Building on the need to assess LLMs in real-world tasks, DOCBench [8] benchmarks document reading systems from domains like academia, finance, and law using an LLM-based evaluation approach. The binary score assignment is done by GPT-4 model to evaluate the correctness of the executed task. While much of the question generation is automated, manual intervention is required for complex question types, limiting scalability. Expanding on specialized task benchmarking, PM-LLM-Benchmark [9] focuses on evaluating LLMs in process mining tasks, addressing the gap in assessing how well these models handle domain-specific knowledge. The benchmark evaluates open-source and commercial LLMs on tasks like process discovery, conformance checking, and root cause analysis. It uses two key implementation paradigms: direct insight provision and code generation. The evaluation relies on an LLM-as-a-Judge system, using GPT-4 to score outputs from 1 to 10. While the mentioned studies are effective and reduce human intervention, reliance on LLMs for evaluation can introduce bias, as LLMs may make subjective judgments.
Some studies have transitioned to automated evaluations by relying on explicit metrics instead of using LLMs as evaluators. CEBench [10] is an example of a fully automated evaluation toolkit, focusing on cost-effectiveness analysis of LLMs using metrics such as response time, memory efficiency, and mean absolute error (MAE). Its customizable setup helps optimise LLM pipelines for both effectiveness and budget and can be used to balance performance with financial considerations. Similarly, CoDI-Eval [11] explores the controllability of LLMs by assessing their ability to execute complex and varied instructions across key tasks such as sentiment, topic, length, keyword usage, and toxicity avoidance. The benchmark highlights the challenges LLMs face when adhering to explicit constraints and aims to improve alignment with human expectations. By integrating a broad range of instructions, CoDI-Eval evaluates both open-source and commercial models, including GPT-4 and Llama, using accuracy-based metrics to measure adherence to instructions without relying on subjective judgments or LLMs for scoring. These benchmarks demonstrate the potential to reduce human reliance while maintaining robust, measurable outcomes through explicit evaluation frameworks; however, they require specifically designed metrics tailored to different contexts for accurate assessments.
2.2 Existing Studies on Persona-aware Response Generation
Personalization, emerging as a crucial aspect of LLM usage, has been addressed through various evaluations, including published benchmarks. Focusing on released benchmarking, the release of RPBench-Auto by BosonAI [2] highlights this evaluation gap. The benchmark is built on persona-augmented conversations sourced from online communities. Inspired by ArenaHard [12] and AlpacaEval [13], RPBench-Auto uses GPT-4-Turbo as a judge model to compare responses from evaluated models against a GPT-4o baseline in character-based and scene-based role-plays. Although effective, this approach is prone to biases inherent in both the baseline and judge models. Their evaluation relies on a win rate metric over a limited test set (80 personae), rather than explicit evaluation metrics, to measure how often the evaluated model’s responses are preferred over the baseline. The study underscores the pressing need for a specialized benchmark in personalized response generation, and an analysis of the CoT effect could further enhance their setup. In the same context, TIMECHARA [3] benchmarks the temporal consistency of character personas in narratives and the potential hallucinations that can occur in LLMs, focusing on how well these models maintain character integrity over time. The evaluation framework introduces metrics like AlignScore to assess spatiotemporal consistency and a 1-7 Likert scale [14] to measure personality consistency. These evaluations are conducted across various setups, including zero-shot, few-shot, CoT reasoning, and retrieval-augmented generation (RAG). However, the use of LLMs as judges introduces potential biases and increases costs, limiting the feasibility of large-scale assessments. In a related effort, RoleLLM [4] explores a more fine-tuned role-conditioning process to benchmark models’ ability to imitate characters with specific traits. Both open-source and closed-source models are evaluated using a custom dataset, RoleBench. The evaluation incorporates Rouge-L for role-playing accuracy and GPT-4-based human-like evaluations to rank models on their persona-specific performance. While effective at eliciting character-level responses, RoleLLM faces limitations such as reliance on single-turn interactions and manual intervention in data generation, which may affect scalability and the richness of dialogue.
Figure 1: Overview of the PersoBench automatic personalization benchmark framework.
In addition to benchmarking, recent evaluation studies using datasets such as IT-ConvAI2, Blended Skill Talk, and FoCus—also integrated into our work—have focused on assessing dialogue models. For example, the research by Chen et al. (2023) [15] focuses on metrics like BLEU and ROUGE, which, while informative, fail to fully capture the personalization capabilities of more advanced models. Innovative approaches, such as PAL [16] and robust personalized dialogue generation [17], have pushed the field forward, yet the need for more comprehensive evaluation metrics remains evident.
However, the mentioned studies often fall short due to their reliance on a limited array of metrics and older models like GPT-2. Additionally, these studies are constrained by their narrow scope and evaluation frameworks. In contrast, our work comprehensively evaluates both open-source and closed-source LLMs, employing dedicated metrics to assess personalization, fluency, diversity, and coherence. By analyzing model performance in both vanilla and CoT setups within a zero-shot setting, we provide an in-depth comparison across these aspects, setting a new benchmark for evaluating response personalization in LLMs.
@@ 3.1 Problem statement
For our benchmark on personalized response generation, we have adopted the same problem statement as defined by Chen et al. (2023) [17]. Given the dialogue context $C=\{u_{1},\dots,u_{m}\}$ and a set of persona descriptions $P=\{p_{1},\dots,p_{n}\}$, the goal is to generate a personalized response $r$. Formally, the generation problem can be formulated as the following chain rule:
@@
where $\theta$ is the parameters of the dialogue model.
@@ 3.2 Overview of PersonBench
Our automatic benchmarking framework is shown in Fig.1. PersoBench includes around 3,600 samples from three persona-aware datasets, each differing in size and context. It applies eight metrics across four evaluation dimensions based on the datasets and metrics discussed in the recent review study [1]. The framework evaluates the performance of six prominent LLMs—three open-source and three closed-source— within a zero-shot setting under both vanilla and COT prompting setups.
@@ 3.3 Prompt Development
Table 1: Blueprint of the designed prompt for personalized response generation task. The orange text will only be included in the COT setup.
The overall structure of our designed prompt is shown in Table 1. As suggested by related studies [18, 19, 20, 5], the prompt begins with a general problem description to establish context for the LLM, followed by a clear specification of the LLM’s role in addressing the target task. Next, a clear definition of the input is provided, which includes participants’ personas and conversation history, ensuring the necessary information is available for the task. In line with the principles outlined in the literature [21, 22, 18, 23], detailed task instructions are given to guide response generation and mitigate ambiguities. Lastly, the prompt specifies the desired output format for the LLM, supporting both vanilla and COT setups. Specifying the JSON format is crucial for our benchmarking setup, allowing seamless extraction of reasoning and responses for post-processing and evaluation. This strategy is also commonly applied when the output requires more than a simple response [24, 25]. In scenarios where the conversation context involves IT-related queries, our experiments have shown that instructing the LLM to provide a ’natural language response’ is crucial. This prevents the model from generating code instead of a straightforward conversational answer. Based on findings that placing the reasoning before the response improves coherence and adherence to task requirements [24], we adopted this approach in the design of the output structure. To ensure the reasoning does not consume all available tokens and leave insufficient space for the actual response, we included the keywords ’briefly’ and ’(in less than 110)’ in the reasoning instruction. The 110-token limit was determined based on the longest actual response across all selected datasets (shown in Table 2), ensuring an equal token allocation for both the reasoning and the response.
@@ 4 Experiment
4.1 Dataset
To establish a solid foundation for our benchmark, we selected widely recognized datasets in the domain, guided by a recent survey in the same context [1]. Our selection criteria included choosing datasets with a reasonable test/validation set size, which was important given our budget constraints. We prioritized datasets that clearly separated the actual response from the conversation context to meet our data structure requirements and those that required minimal preprocessing to preserve the integrity of their content. Based on these criteria, the following datasets were chosen for our benchmarking experiment, and their statistics are provided in Table 2.
Blended Skill Talk (BST): The BST dataset [^1] https://www.kaggle.com/datasets/thedevastator/multi-modal-conversation-data?resource=download evaluates key conversational competencies, including dialogue flow, topic control, coherence, and the integration of personality, empathy, and knowledge. This dataset is effective for assessing conversational styles, enhancing user engagement, and benchmarking distributed dialogue systems across various modalities.
Follow-up Customized conversation (FoCus) [^2] https://github.com/pkchat-focus/FoCus?tab=readme-ov-file This dataset is designed to generate personalized and knowledgeable responses by grounding them in both the user’s persona and relevant Wikipedia knowledge. Focused on conversations about geographical landmarks, it evaluates how well models integrate persona and knowledge to create customized, contextually appropriate responses, making it valuable for developing advanced conversational agents.
IT-ConvAI2: [^3] https://github.com/CCIIPLab/Persona_Extend/tree/main/datasets/IT_ConvAI2 As an extension of the original ConvAI2 dataset, it is tailored specifically for IT-related conversations. It focuses on dialogues in technical support scenarios, where the conversational agent is expected to assist users with IT issues. As a persona-driven dataset, it includes personalized information about the participants to make the interactions more relevant and context-aware.
Table 2: Characteristics of the selected datasets for benchmarking. ’Resp. Len.’ indicates the overall golden response length range in words.
Name Test Size Context Type Query Sep U/D Role Resp. Len.
@@ 4.1 Dataset
Table 3: Fluency evaluation of LLMs on the IT-ConvAI2 dataset.
Model Vanilla Setup COT Setup
ROUGE (R1/R2/RL) METEOR BERTScore (Prec/Rec/F1) ROUGE (R1/R2/RL) METEOR BERTScore (Prec/Rec/F1)
Mistral 7B 0.12/0.02/0.11 0.14 0.81/0.81/0.81 0.09/0.01/0.09 0.11 0.80/0.80/0.80
Qwen2 7B 0.10/0.02/0.10 0.14 0.83/0.84/0.83 0.09/0.01/0.08 0.12 0.82/0.83/0.83
Llama3.1 8B 0.11/0.02/0.10 0.13 0.76/0.77/0.76 0.11/0.02/0.10 0.12 0.84/0.83/0.84
GPT3.5 Turbo 0.12/0.02/0.11 0.15 0.85/0.85/0.85 0.11/0.02/0.10 0.13 0.84/0.84/0.84
GPT4o Mini 0.11/0.02/0.10 0.16 0.84/0.85/0.84 0.10/0.01/0.10 0.16 0.83/0.85/0.84
GPT4 Turbo 0.12/0.02/0.11 0.15 0.85/0.86/0.85 0.11/0.02/0.10 0.13 0.84/0.84/0.84
Table 4: Fluency evaluation of LLMs on the Blended Skill Talk dataset.
Model Vanilla Setup COT Setup
ROUGE (R1/R2/RL) METEOR BERTScore (Prec/Rec/F1) ROUGE (R1/R2/RL) METEOR BERTScore (Prec/Rec/F1)
Mistral 7B 0.10/0.01/0.10 0.11 0.80/0.81/0.81 0.10/0.01/0.09 0.11 0.80/0.82/0.81
Qwen2 7B 0.11/0.01/0.10 0.12 0.82/0.84/0.83 0.09/0.01/0.09 0.11 0.83/0.86/0.84
Llama3.1 8B 0.12/0.02/0.11 0.12 0.81/0.82/0.82 0.11/0.01/0.10 0.11 0.85/0.86/0.85
GPT3.5 Turbo 0.13/0.02/0.11 0.13 0.85/0.86/0.86 0.11/0.01/0.10 0.12 0.82/0.84/0.83
GPT4o Mini 0.13/0.02/0.12 0.14 0.85/0.87/0.86 0.12/0.01/0.11 0.14 0.84/0.86/0.85
GPT4 Turbo 0.12/0.02/0.11 0.12 0.85/0.86/0.86 0.11/0.01/0.10 0.12 0.84/0.86/0.85

### v2

@@ Abstract
While large language models (LLMs) have exhibited impressive conversational capabilities, their proficiency in delivering personalized responses remains unclear. Although recent benchmarks automatically evaluate persona consistency in role-playing contexts using LLM-based judgment, the evaluation of personalization in response generation remains underexplored. To address this gap, we present an automated benchmarking pipeline, PersoBench, to evaluate the personalization ability of LLMs in persona-aware dialogue generation within a zero-shot setting. Our framework employs a structured pipeline comprising speaker-aware annotation, task-specific and context-driven prompt construction, response post-processing, and automated evaluation across multiple dimensions of generation quality. In particular, the pipeline performs text preprocessing and speaker labeling, constructs structured prompts with task instructions and LLM roles, validates response format, and evaluates valid outputs across fluency, personalization, diversity, and coherence. We assess the performance of four open-source and four closed-source LLMs using well-known datasets and a range of explicit metrics. Our findings reveal that while LLMs excel at generating fluent and diverse responses, they are far from satisfactory in delivering personalized and coherent responses, considering both the conversation context and the provided personas.
Keywords Personalized Response Generation, Conversational AI, Natural Language Processing, Large Language Models, Benchmarking
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized NLP, excelling in human-like text generation across domains and becoming central to dialogue systems. Yet, as LLMs are increasingly deployed in real-world applications, such as customer support [5], mental health counseling [17, 22], and educational tutoring [36], the ability to generate personalized responses is no longer optional. Personalization enhances user satisfaction, trust, and task success by adapting responses to individual preferences, goals, and conversational history [31]. Without personalization, even fluent and coherent outputs risk feeling generic or misaligned with user intent. Despite this growing demand, evaluating LLMs’ capacity for personalization remains underexplored. Existing benchmarks focus on role adherence or character simulation, such as RPBench-Auto [2], TIMECHARA [1], and RoleLLM [40], but none offer a dedicated framework for assessing automatic personalized response generation. Moreover, these benchmarks often rely on LLMs as evaluators, introducing bias, and suffer from limited experimental scope [15, 41, 2]. This paper addresses these gaps by proposing a benchmarking pipeline specifically designed to evaluate personalized response generation, with robust evaluation protocols and diverse user contexts.
To fill this gap, we introduce PersoBench, an automated benchmarking pipeline for response personalization, to assess the strengths and limitations of current LLMs in generating personalized responses. To the best of our knowledge, no prior work has introduced a comprehensive pipeline specifically focused on evaluating response personalization in LLMs. Using comprehensive datasets and a diverse set of established metrics, including fluency, diversity, coherence, and personalization, we ensure a robust evaluation of various aspects of response generation, drawing on insights from a recent survey in the field [10]. While fluency and diversity assess the fundamental requirements of a dialogue system, coherence and personalization evaluate response alignment with the provided user persona. More specifically, in line with this objective of the mentioned context, we aim to answer the following research questions:
@@
    To what extent can LLMs generate natural responses that are linguistically fluent?
@@
    How effectively can LLMs generate personalized responses that reflect the provided persona?
@@
    To what extent do LLMs generate coherent responses aligned with the dialogue context and persona?
To answer the above questions, we benchmarked four open-source and four closed-source LLMs across three popular persona-aware conversational datasets, each highlighting a distinct facet of personalization such as open-domain interests, structured traits, or task-specific roles. This multi-perspective design enables a more robust and generalizable evaluation of personalized response generation. The LLMs’ performance are evaluated in a zero-shot setting to benchmark their inherent capabilities without introducing examples that could enhance their performance, ensuring the results reflect their baseline ability. This approach is applied to both standard and Chain of Thought (CoT) reasoning configurations. Our evaluation goes beyond assessing overall response quality and contextualization, employing eight established metrics in conversational AI to gauge performance across fluency, diversity, coherence, and personalization dimensions. This work contributes by (1) offering a multi-aspect evaluation of LLM personalization in zero-shot settings, (2) investigating the influence of CoT prompting on response personalization, and (3) assessing response time and instructability (the ability to adhere to task instructions) based on LLMs’ adherence to task instructions.
@@
The implementation of PersoBench, including evaluation scripts and results, is publicly available at: github.com/salehafzoon/PersoBench.
@@ 2 Related Work
2.1 Persona-aware Response Generation
Personalized response generation has emerged as a key challenge in open-domain dialogue systems, where maintaining consistency with user-specific traits is essential for engagement and realism [27]. A range of modeling strategies has been proposed to address this challenge, focusing on persona conditioning, stylistic control, and long-term coherence.
To mitigate temporal drift in character-driven interactions, TIMECHARA [1] is introduced to model time-sensitive persona attributes. By encoding temporal markers and character evolution, this approach improved consistency across multi-turn dialogues. RoleLLM [40] extended this idea by incorporating RoleBench and RoCIT, which enabled role-conditioned generation through structured prompts and contextual role embeddings. Evaluation logic was embedded within the generation pipeline, allowing for internal consistency checks. Stylistic personalization was explored in RAGs to Style [33], where retrieval-augmented generation was combined with style embeddings to guide output fluency and tone. This method retrieved relevant persona-aligned examples and fused them with stylistic constraints during decoding. In contrast, PAPER [25] adopted a CoT framework to interpret and expand persona tags prior to generation, enhancing transparency and interpretability. PPlug [28] proposed a plug-and-play module that attaches user embeddings to LLM inputs, enabling scalable personalization without retraining or fine-tuning.
These approaches reflect a spectrum of personalization strategies, including prompt engineering, retrieval-based conditioning, modular embedding, and reasoning over persona traits. Despite their innovation, evaluation remains fragmented. Metrics such as Persona-F1 [20], AlignScore [42], and win-rate have been commonly employed, yet they lack standardization and often fail to capture multi-dimensional traits such as coherence and instructability.
2.2 LLM Evaluation Approaches
The evaluation of LLMs has traditionally relied on benchmark datasets that offer standardized corpora and annotated references across diverse tasks. Foundational suites such as GLUE [39] and SuperGLUE [38] were introduced to assess general language understanding through tasks like sentiment classification, natural language inference, and commonsense reasoning. These benchmarks were later extended by MMLU [18], which evaluated LLMs across 57 academic subjects, including law, medicine, and mathematics, under zero-shot and few-shot settings. More specialized datasets have also been proposed to target domain-specific capabilities. For instance, FELM [44] was designed to assess factuality across five knowledge domains using curated reference links and segment-level annotations, while DOCBench [46] compiled real-world documents and structured queries to evaluate document comprehension. PM-LLM [4] focused on process mining tasks, and TrustGPT [19] introduced ethically sensitive prompts to measure bias, toxicity, and value alignment.
Alongside these datasets, explicit metrics have been widely adopted to quantify model output quality. ROUGE [26], and BERTScore [43] have served as standard tools for measuring lexical overlap and semantic similarity between generated and reference texts. Despite their computational efficiency, these metrics have been limited in capturing deeper qualities such as coherence, instructability, and persona fidelity, especially in open-ended or stylistically diverse generation tasks.
To address these limitations, human evaluation has been increasingly employed to assess subjective dimensions of model behavior. In TrustGPT [19], human annotators were used to validate toxicity and bias judgments across multiple alignment tasks, while FELM [44] incorporated segment-level human ratings to benchmark factual consistency. These efforts have underscored the importance of human judgment in capturing nuanced traits that automatic metrics often overlook.
More recently, model-based evaluation techniques have gained traction. The LLM-as-a-Judge paradigm [16] has been widely adopted, wherein a powerful model (e.g., GPT-4 or Claude) is prompted to assess the output of another model across multiple dimensions such as helpfulness, factuality, and stylistic alignment. This approach enables scalable and flexible evaluation, often outperforming traditional metrics in capturing coherence and relevance. However, concerns have been raised regarding reproducibility, evaluator bias, and architectural entanglement, particularly when the evaluator and evaluated models share similar design principles.
Figure 1: Overview of the PersoBench automatic personalization benchmarking pipeline.
2.3 LLM Evaluation Frameworks
While many evaluation approaches for LLMs have focused on task-specific datasets and isolated metrics, a growing body of work has sought to formalize reusable frameworks that support structured, multi-dimensional assessment. These systems aim to standardize evaluation pipelines, automate scoring, and improve reproducibility across diverse model architectures and tasks.
Early efforts emphasized general-purpose evaluation toolkits. DeepEval [^1] https://github.com/confident-ai/deepeval introduced a unit-test-inspired framework that supports over 14 metrics, including contextual relevance, hallucination detection, and bias analysis. Its integration with Pytest and synthetic dataset generation has made it suitable for production-level monitoring. MLFlow’s LLM Evaluate module [^2] https://mlflow.org/docs/latest/llm-evaluate/index.html provided lightweight support for retrieval-augmented generation and question answering pipelines, emphasizing developer-friendly integration. Deepchecks [^3] https://www.deepchecks.com/llm-evaluation-framework-steps-components/ extended this paradigm by combining LLM-as-a-Judge with human-in-the-loop review, offering trace tagging and safety diagnostics for enterprise deployment.
More comprehensive frameworks have been proposed to evaluate controllability, coherence, and factuality. CoDI-Eval [11] introduced a diversified instruction evaluation protocol, enabling structured assessment of controllability across open-ended prompts. UniEval [45] offered a general-purpose evaluation framework based on fine-tuned T5 models, supporting dimensions such as relevance, fluency, consistency, and factuality. Its modular design and task-agnostic scoring have made it widely applicable in summarization, dialogue, and QA tasks. CEBench [37], while primarily focused on cost-effectiveness, proposed a configuration-based evaluation toolkit that jointly considers latency, throughput, and token usage, highlighting trade-offs between performance and deployment feasibility.
Frameworks tailored to persona-aware evaluation have also been proposed. RPBench-Auto [2] introduced an automated evaluation pipeline using GPT-4-Turbo as a judger model, assessing persona consistency, fluency, and alignment through structured prompts and scoring templates. RoleLLM [40], though primarily a generation model, embedded evaluation logic via RoleBench and RoCIT to assess role fidelity and contextual coherence. These systems reflect a growing emphasis on personalization-aware evaluation, where traits such as instructability, style alignment, and long-term persona fidelity are jointly assessed.
@@ 3.1 Problem statement
For our benchmarking pipeline on personalized response generation, we have adopted the same problem statement as defined by Chen et al. (2023) [9]. Given the dialogue context $C=\{u_{1},\dots,u_{m}\}$ and a set of persona descriptions $P=\{p_{1},\dots,p_{n}\}$, the goal is to generate a personalized response $r$. Formally, the generation problem can be formulated as the following chain rule:
@@
where $\theta$ is the parameter of the dialogue model and $r_{t}$ represents the response generated for the given context $C$ at step, which refers to the most recent utterance in the dialogue.
Unlike role-playing, which simulates specific characters with detailed profiles and temporal constraints, personalization focuses on generating coherent, contextually relevant responses aligned with user-defined personas, without requiring rich backstories.
@@ 3.2 Overview of PersonBench
Our automatic benchmarking framework is shown in Fig.1. PersoBench includes around 3,600 samples from three persona-aware datasets, each differing in size and context. It applies eight metrics across four evaluation dimensions based on the datasets and metrics discussed in the recent review study [10]. The framework evaluates the performance of eight prominent LLMs, four open-source and four closed-source, within a zero-shot setting under both vanilla and CoT prompting, where LLMs are instructed to provide reasoning as well.
@@ 3.3 Prompt Development
Table 1: Prompt blueprint for personalized response generation. Orange text appears only in the CoT setup.
@@
The overall structure of our designed prompt is shown in Table 1. As suggested by related studies [35, 8, 30, 44], the prompt begins with a general problem description to establish context for the LLM, followed by a clear specification of the LLM’s role in addressing the target task. Next, a clear definition of the input is provided, which includes participants’ personas and conversation history, ensuring the necessary information is available for the task. In line with the principles outlined in the literature [6, 34, 35, 3], detailed task instructions are given to guide response generation and mitigate ambiguities. Lastly, the prompt specifies the desired output format for the LLM, supporting both vanilla and CoT setups. Specifying the JSON format is crucial for our benchmarking setup, allowing seamless extraction of reasoning and responses for post-processing and evaluation. This strategy is also commonly applied when the output requires more than a simple response [14, 21]. We adopt a zero-shot setting to evaluate LLMs, ensuring the results reflect their baseline performance. This approach also prevents masking any limitations by relying on external examples or prompt engineering, which would be misaligned with our objective of analyzing the models’ true generalization capabilities. In IT-related queries, instructing the LLM to provide a ’natural language response’ is crucial to avoid generating code instead of conversational answers. Following findings that reasoning before the response improves coherence and task adherence [14], we adopted this structure. To ensure reasoning remains concise and leaves space for the response, we added keywords like ’briefly’ and ’(in less than 110)’ in the instruction. The 110-token limit, derived from the longest response in our datasets (Table 2), ensures balanced token allocation.
@@ 4 Experiment
To establish a solid foundation for our benchmarking pipeline, we selected widely recognized datasets in the personalized dialogue domain, guided by a recent survey [10]. Practical constraints such as test/validation set availability and the presence of clearly separated response fields guided our selection. Importantly, we sought datasets that require minimal preprocessing and exhibit diverse personalization properties, enabling a more comprehensive evaluation. While each dataset has been used in prior studies individually, they differ significantly in the aspects of personalization they emphasize. By bringing them together under a shared evaluation protocol, we enable holistic, cross-dataset benchmarking of large language models. The following datasets were selected, and their statistics are reported in Table 2:
Each dataset captures a different facet of personalization, from light-touch social engagement (BST) to semi-structured interest modeling (FoCus) and domain-specific technical traits (IT-ConvAI2). While prior work has examined them separately, our benchmarking pipeline unifies these resources, enabling evaluation across diverse personalization regimes. This integration yields more generalizable insights into how models adapt to varied user traits, conversation types, and task settings, offering a broader and more realistic view of personalized generation performance.
Table 2: Characteristics of selected datasets for benchmarking. ’R. Len’ stands for the golden response length range in words, and ’Size’ denotes the test size.
Name Size Context Type Query Sep U/D Role R.Len
@@
Table 3: Fluency evaluation across the datasets in Vanilla and CoT setups. B-F1 refers to BERTScore-F1, and NUE refers to the Naturalness evaluation metric within the UniEval framework.
Model IT-ConvAI2 BST FoCus
Vanilla CoT Vanilla CoT Vanilla CoT
B-F1 NUE B-F1 NUE B-F1 NUE B-F1 NUE B-F1 NUE B-F1 NUE
Mistral 7B 0.81 0.89 0.80 0.90 0.81 0.90 0.81 0.90 0.48 0.53 0.76 0.82
Qwen2 7B 0.83 0.94 0.83 0.93 0.83 0.93 0.84 0.93 0.42 0.46 0.62 0.63
Gemma 7B 0.83 0.92 0.75 0.85 0.83 0.93 0.79 0.87 0.49 0.54 0.72 0.77
Llama3.1 8B 0.76 0.85 0.84 0.94 0.82 0.92 0.85 0.95 0.47 0.51 0.80 0.85
Gemini 1.5 Pro 0.80 0.90 0.84 0.94 0.83 0.92 0.84 0.92 0.66 0.70 0.71 0.76
GPT3.5 Turbo 0.85 0.96 0.84 0.96 0.86 0.96 0.83 0.92 0.78 0.84 0.80 0.86
GPT4o Mini 0.84 0.95 0.84 0.95 0.86 0.96 0.85 0.95 0.80 0.85 0.86 0.93
GPT4 Turbo 0.85 0.96 0.84 0.96 0.86 0.96 0.85 0.96 0.84 0.91 0.86 0.93